Status Report

NASA ISS On-Orbit Status 3 October 2008

By SpaceRef Editor
October 3, 2008
Filed under , , ,
NASA ISS On-Orbit Status 3 October 2008

All ISS systems continue to function nominally, except those noted previously or below.

FE-2 Chamitoff started out in the JAXA Kibo JPM (JEM Pressurized Module) by preparing the AVCO (Air Ventilation Closeout) for the subsequent MELFI rack transfer. [The reconfiguration consisted in removing a soft dummy panel from location A4 and relocating a hard dummy panel from the D4 bay to A4, torquing it down and fastening its ground strap to a standoff.]

Assisted by CDR Volkov where necessary, the FE-2 then relocated the MELFI (Minus-Eighty Laboratory Freezer for ISS) rack from the US Lab (O4 bay) to the JPM (D4 bay). [For the long-planned transfer to the Japanese Kibo, Gregory had to disconnect/reconnect umbilicals, remove fasteners, release capture mechanisms, disassemble braces, free the rack from its position, “fly” it to the JPM, install it at D4, perform post-transfer checkouts on the MELFI electronics box and PCS (Portable Computer System) laptop and reconfigure it and the rack for future use. Afterwards, temporarily removed items were to be restowed and everything cleaned up. Unlike the other racks, during the transfer MELFI was on a “thermal clock”, viz.: once its internal Brayton motor (for cooling) was turned off, there was an 8-hr window to get the freezer rack turned on again. If problems cropped up during the relocation (e.g., with a QD/quick disconnect) which would have exceeded the window, the crew had a back-out plan to return the rack to LAB1O4 for reconnection.]

Then, the FE-2 worked on the JPM’s TCA (Thermal Control Assembly) Low Temperature Loop by repositioning its gas trap bypass valve for upcoming gas trap activities.

Also in the Kibo module, Gregory supported ongoing Marangoni Surface experimentation in the FPEF (Fluid Physics Experiment Facility), ground-commanded from the SSIPC (Space Station Integration & Promotion Center) at Tsukuba/Japan, by activating the MMA (Microgravity Measurement Apparatus) and its laptop (MLT), first powering up the MMA’s NCU/RSU (Network Control Unit/Remote Sensor Unit) set from the Ryutai rack’s UDC (Utility DC-to-DC Converter), then turning on both NCU/RSU and MLT. [Tsukuba is conducting the Marangoni experiment basically only during crew sleep (for quiescence), starting again tonight until 10/16 (except for 10/9). Crewmembers can enter Kibo during the experiment but are advised not to cause any major disturbance of the micro-G environment.]

Later, Chamitoff updated the three copies of the SODF (Systems Operation Data File) EMER-2 emergency books from the SM (Service Module), FGB and Lab. [The books were updated with new pages on Fire Source Locations in Lab, COL (Columbus Orbital Laboratory) and JPM.]

In the SM, FE-1 Kononenko performed the periodic download of protocol/log files of the BRI Smart Switch Router to the RSS1 laptop for subsequent dumping to the ground via OCA, for specialists to review performance parameters, then rebooted the RRS1. [BRI is part of the RS OpsLAN network with connections to the three SSC clients, the relatively new Ethernet tie-in with the US network, and a network printer in the RS (Russian segment).]

Continuing the current round of the monthly preventive maintenance of RS (Russian segment) ventilation systems, Volkov performed a 1h15m inspection and cleaning of Group A ventilator fans and grilles in the SM, while Kononenko a bit later conducted ventilation system maintenance in the DC1 (Docking Compartment) by cleaning the PF1 & PF2 dust collectors in its air duct system and the protective mesh screens of the V1 & V2 ventilator fans.

Greg Chamitoff took measurements for the regular atmospheric status check for ppCO2 (Carbon Dioxide partial pressure) in the Lab, SM at panel 449 and COL (Columbus Orbital Laboratory) plus battery ticks, using the hand-held CDMK (CO2 Monitoring Kit, #1002). The unit was then deactivated and returned to its stowage location (LAB1S2). [Purpose of the 5-min activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements. Measurements were 0.47% in SM, 0.47% in US Lab, 0.46% in COL.]

The FE-2 performed the periodic (monthly) battery check and reboot of all active US PCS (Portable Computer System) and the COL PWS (Portable Workstation) laptops.

Oleg performed the routine daily servicing of the SOZh system (ECLSS, Environment Control & Life Support Systems) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Later, Kononenko also conducted the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

The CDR & FE-1 had another 4h10m scheduled between them for stowage activities on Progress M-65/30P, going by an uplinked list of items to be discarded and their stowage locations. [30P currently also contains liquid waste (urine), pumped to water tank BV1 from 8 EDV-U containers and to BV2 from 6 EDV-Us. Solid waste is stored in 10 KTO containers, besides discarded BKO & BRPK condensate hardware, 22 food ration containers, 12 used dust collector cartridges, a replaceable SPN pump unit, and other items of trash.]

Oleg & Sergey each had the daily one-hour period to themselves for the regular crew departure preparations, working on the standard end-of-increment cleanup preparatory to their return to Earth later this month. [It is usual for Russian crewmembers to be granted reduced workdays for making their departure preparations, as their return date approaches.]

The crew completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR/2.5h, FE-1/2.5h), and RED resistive exercise device (FE-2).

Later, the FE-2 is to transfer the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Afterwards, Gregory will be starting his first SOLO (Sodium Loading in Microgravity) session, which runs for five days. [During this period, Chamitoff follows a special high-salt diet, for which prepared meals are provided onboard. All three daily meals will be logged on sheets stowed in the PCBA (Portable Clinical Blood Analyzer) Consumable Kit in the MELFI along with control solution and cartridges for the PCBA. SOLO, an ESA/German experiment from the DLR Institute of Aerospace Medicine in Cologne/Germany, investigates the mechanisms of fluid and salt retention in the body during long-duration space flight. Background: The hypothesis of an increased urine flow as the main cause for body mass decrease has been questioned in several recently flown missions. Data from the US SLS1/2 missions as well as the European/Russian Euromir `94 & MIR 97 missions show that urine flow and total body fluid remain unchanged when isocaloric energy intake is achieved. However, in two astronauts during these missions the renin-angiotensin system was considerably activated while plasma ANP concentrations were decreased. Calculation of daily sodium balances during a 15-day experiment of the MIR 97 mission (by subtracting sodium excretion from sodium intake) showed an astonishing result: the astronaut retained on average 50 mmol sodium daily in space compared to balanced sodium in the control experiment.]

At ~4:15am EDT, the crew held the regular (nominally weekly) tagup with the Russian Flight Control Team (GOGU), including Shift Flight Director (SRP), at TsUP via S-band/audio, phone-patched from Houston and Moscow.

At ~4:30am, the two cosmonauts linked up with TsUP-Moscow stowage specialists via S-band to conduct the weekly IMS tagup, discussing stowage issues and equipment locations.

At ~11:00am, Sergey, Oleg and Greg supported two live PAO TV interviews with US clients,- CBS News (Bill Harwood), and KHOU-TV “Great Day, Houston” (Deborah Duncan).

At ~3:55pm, the crewmembers will convene for their standard bi-weekly teleconference with the JSC Astronaut Office (Steve Lindsey), via S-band S/G-2 audio & phone patch.

As generally every day now, today starting at ~9:00am and running until 3:00pm, the US CDRA (Carbon Dioxide Removal Assembly) is being activated intermittently for two half-cycles to control ppCO2 levels. This configuration for the daily ops does not require connecting & disconnecting the ITCS cooling loop. [A forward plan is in work for cycling the CSV (CO2 Selector Valve) to prevent its sticking. CDRA remains “yellow” on the ISS critical systems list.]

WRM Update: An updated WRM (Water Recovery Management) “cue card” was uplinked overnight for the crew’s reference, updated with yesterday’s water audit. [The new card (17-1016B) lists 29 CWCs (~962.3 L total) for the four types of water identified on board: technical water (283.2 L, for Elektron electrolysis, except for 22.2 L off-limit because of Wautersia bacteria), potable water (627.6 L, incl. 174.6 L currently on hold), condensate water (34.5 L), waste/EMU dump and other (17 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

New tasks currently on Greg Chamitoff’s discretionary US “job jar” task list are –

  • Labeling a new PMIC (Portable Microphone) & HHM (Hand-Held Microphone),
  • Hard-wiring & reconfiguring the SSC-15 (Station Support Computer 15) laptop in the Kibo JPM for wireless operation via the new WAPs (Wireless Access Points), thus getting rid of the drag-through Ethernet cable (which could slow down rapid hatch closure); and
  • Searching for MSG (Microgravity Science Glovebox) glove rings, then dividing them into old vs. new.

Reboost Planning: Planning for tomorrow’s ISS reboost is underway, with details TBD. Attitude control authority handover to RS thrusters, for maneuvering to burn attitude and conduct the firing, is expected for 7:30am EDT, with return to US Momentum Management at ~8:30am.

Week 24 Scheduled Main Activities:

  • Sat. (10/4): Station cleaning; SAMS activation; SOLO Diet monitor; ISS reboost (~7:45am EDT);
  • Sun. (10/5): PFP-ODNT OBT/trng.; SODF updating.

CEO (Crew Earth Observations) photo targets uplinked for today were Karakoram (this target area, NW of the Himalaya, is considered as one of the Greater Ranges of Asia. With over 60 peaks above 7,000m, it is also the most heavily glaciated area of the world outside the Polar Regions. ISS had a fair-weather pass before the start of the snow season in early afternoon sun. Looking left of track and trying for context views of this beautiful area using the short lens settings), Hurricane Ike Devastation (DYNAMIC EVENT: There is ongoing interest in views of the extent and impact of the recent devastation of the upper Texas coast by Hurricane Ike. Greg’s early afternoon pass should have offered clear, oblique views of SE Texas as he tracked NE-ward over the Gulf, parallel to the coast. Using the long lens settings and looking left of track for extensive debris fields and possible coastline changes from the station’s unique vantage point), Madrean Sky Islands (these ecologically diverse “islands” exist in the higher elevations of the mountain ranges of the Sonora desert of northwestern Mexico and the southwestern US. As ISS tracked NE-ward over the Baja Peninsula, Greg was to shoot right of track for oblique contextual views of this region using the short lens settings. Fair weather was anticipated), and Coral reefs, American Samoa (although much of this target area was to the right of track, Greg appeared to have an excellent late morning pass with only a few clouds. It should have provided nadir views of the islands of Western Samoa and Apia. Using the long lens settings and trying to map in detail the beautiful fringing reef structures of these islands).

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 8:45am EDT [= epoch]):
Mean altitude — 352.0 km
Apogee height — 356.2 km
Perigee height — 347.7 km
Period — 91.58 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006307
Solar Beta Angle — -8.2 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.72
Mean altitude loss in the last 48 hours — 73 m
Revolutions since FGB/Zarya launch (Nov. 98) — 56558

Significant Events Ahead (all dates Eastern Time, some changes possible.):
10/04/08 — ISS Reboost (~7:45am)
10/12/08 — Soyuz TMA-13/17S launch (~3:03am EDT; Fincke, Lonchakov, Garriott)
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port, ~4:51am)
10/24/08 — Soyuz TMA-12/16S undocking (DC1 nadir) & landing
11/02/08 — Progress 30P reboost; Daylight Saving Time (DST) ends.
11/16/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC (~7:02pm EST) – U/R 11/14?
11/18/08 — STS-126/Endeavour/ULF2 docking – U/R
11/20/08 — ISS 10 Years
11/25/08 — Progress M-65/30P undocking & deorbit
11/26/08 — Progress M-66/31P launch
11/30/08 — Progress M-66/31P docking
12/01/08 — STS-126/Endeavour/ULF2 landing (~1:25pm EST est.)
02/09/09 — Progress M-66/31P undocking & deorbit
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 – Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress M-67/32P undocking & deorbit
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/25/09 — Soyuz TMA-15/19S launch
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 — STS-128/Atlantis/17A – MPLM(P), last crew rotation
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.